OpenStudy (anonymous):

Jill sold half of her comic books and then bought eleven more. She now has 28. With how many did she begin? WRITE THE EQUATION

OpenStudy (stormfire1):

28 = 1/2x + 11 28-11 = 1/2x 34 = x

OpenStudy (anonymous):

Let 'X' be the no.of comics Jill had initially. She sold half = X/2 She bought 11 more = X/2+11 Now she has 28 which means her total collection is now 28 X/2+11 = 28 => X/2=17 => X=34

OpenStudy (anonymous):

(1/2)x + 11=28 where x represents the number of comic books; subtract 11 =>(1/2)x = 17; multiply by 2 => x=34